OpenStudy (anonymous):

Solve the polynomial function. Find all of the zeros; there are five zeros. Here is the equation: X^5 - 8x^4 + 28x^3 - 56x^2 +64x - 32 Use synthetic division/quadratic equation :) Solve the polynomial function. Find all of the zeros; there are five zeros. Here is the equation: X^5 - 8x^4 + 28x^3 - 56x^2 +64x - 32 Use synthetic division/quadratic equation :) @Mathematics

OpenStudy (anonymous):

The rational roots are +/- divisors of 32. I will give you hint that 2 is a root. Divide your polynomial by x-1 to get \[ \frac{x^5-8 x^4+28 x^3-56 x^2+64 x-32}{x-2}=x^4-6 x^3+16 x^2-24 x+16 \] Try to find some roots of \[x^4-6 x^3+16 x^2-24 x+16 \]

OpenStudy (anonymous):

Also x=2 is a root of \[x^4-6 x^3+16 x^2-24 x+16=0 \] then \[ \frac{x^4-6 x^3+16 x^2-24 x+16}{x-2}=x^3-4 x^2+8 x-8 \]

OpenStudy (anonymous):

Aagain x=2 is a root of \[x^3-4 x^2+8 x-8=0 \] So \[ \frac{x^5-8 x^4+28 x^3-56 x^2+64 x-32}{(x-2)^3}=x^2-2 x+4\\ x^5-8 x^4+28 x^3-56 x^2+64 x-32=(x-2)^3 \left(x^2-2 x+4\right) \] Use the the quadratic formula for \[x^2-2 x+4=0 \]

OpenStudy (anonymous):

You get \[\begin{array}{c} x=1-i \sqrt{3} \\ x=1+i \sqrt{3} \end{array} \] in addtion you have 2 as a triple root